The ratio of the length and breadth of a plot is 4 : 5. I f the length is 20 m less than the breadth, what
is the perimeter of the plot?

Answer by checkley79(3341)

4:5=(X-20):X CROSS MULTIPLY.
5(X-20)=4X
5X-100=4X
5X-4X=100
X=100 ANS FOR THE LONGER SIDE.
100-20=80 FOR THE SHORTER SIDE.
PROOF:
4/5=80/100
4/5=4/5
PERIMETER=2L+2W
P=2*80+2*100
P=160+200
P=360m. ANS.